OpenStudy (anonymous):

What are the real or imaginary solutions of each polynomial equation? Show work. x^4-13x^2+36=0 x^3-8=0

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

FActor this: x^4-13x^2+36=0

OpenStudy (anonymous):

I can work with real solutions, but when it comes to imaginary it really messes me up.

OpenStudy (anonymous):

Ok, hold on a sec.

Directrix (directrix):

Factor as a trinomial like y^2 - 13y + 36 = 0 but your y will be x^2

OpenStudy (anonymous):

So, (x^2-9) (x^2-4)?

Directrix (directrix):

Yes, and factor each of the binomials again. They are differences of squares.

OpenStudy (anonymous):

Wait so then would the solutions be + or - 3 and + or - 2?

Directrix (directrix):

Work in steps. It is not time to get to an answer yet. You are supposed to factor each of these: (x^2-9) (x^2-4)

Directrix (directrix):

Factor. That is all at this point.

OpenStudy (anonymous):

So... x^2-9=0 x^2=9 Square each side x=square root of 9 x=+ or - 3 And the other one x^2-4=0 x^2=4 Square each side x=square root of 4 x=+ or - 2

OpenStudy (anonymous):

Oh wait I did that wrong!

OpenStudy (anonymous):

Hold on I think I can fix it.

Directrix (directrix):

No, FACTOR x^4 -13x^2+36 =0 (x^2-9) (x^2-4) = 0 (x + 3)* (x - 3) * ( x + 2) * (x - 2) = 0 Do not lose sight of the entire problem.

Directrix (directrix):

(x^2-9) (x^2-4) = (x + 3)* (x - 3) * ( x + 2) * (x - 2) That is factoring.

OpenStudy (anonymous):

Yeah I just realized that! I'm doing this on paper with me as well so I just saw that. I totally forgot what you meant by factoring it. I'm so sorry. But yes that's what I got as well. And then I'd solve for 0 and then I'd get.. wait, it's still the same answer tho right?

OpenStudy (anonymous):

I'm sorry lol that part just flew past my head.

Directrix (directrix):

It may be but here's the deal. All problems of this sort don't factor so easily and have integer roots. So, it is good to have a procedure. There is no need to be sorry. You are learning. I'm working on paper, too.

OpenStudy (anonymous):

Alright, so it's important that I do that step as well. I'll remember that next time, thanks :)

Directrix (directrix):

Next Step: (x + 3)* (x - 3) * ( x + 2) * (x - 2) = 0 Use the Zero Product Property to get the roots

Directrix (directrix):

x + 3 = 0 OR x-3 = 0 or x +3 = 0 or x -2 = 0

OpenStudy (anonymous):

So set each one equal to 0 and solve correct? x+3=0 x=-3 x-3=0 x=3 x+2=0 x=-2 x-2=0 x=2 So the solutions to this polynomial is x=-3, -2, 2, 3

Directrix (directrix):

Correct.

OpenStudy (anonymous):

Thank you so much! :) But on the other one, after bringing the 8 to the other side and having x^3=8, I have no idea what to do then. Maybe a very vague idea but not a clear one.

Directrix (directrix):

What about this: x^3-8=0 Have you memorized the difference of two cubes factoring formula?

OpenStudy (anonymous):

I think I have it written down here... a^3-b^3=(a-b)(a^2+ab+b^2)

Directrix (directrix):

Here's a chart.

OpenStudy (anonymous):

Wait, could I do x^3-2^3=0, and then factor then solve?

Directrix (directrix):

So apply that to x^3-8 = 0 Factor the x^3 - 8

Directrix (directrix):

Right. The second factor of x^3 - 8 is where you'll see some imaginaries (Complex roots) Just factor and we'll worry about that later.

OpenStudy (anonymous):

Alright, give me a sec :)

Directrix (directrix):

Wait, could I do x^3-2^3=0, and then factor then solve? YES

OpenStudy (anonymous):

Ok here's what I got: x^3-2^3=(x-2)(x^2+2x+8) =(x-2)(x+4)(x+2)

Directrix (directrix):

x^3 - 2^3 = (x-2)* (x^2 + 2x +4) = 0

OpenStudy (anonymous):

Wait why would it be 4 not 8? The chart says y^3. So isn't 2^3 equal to 8?

Directrix (directrix):

Look at the factoring formula. (x-2)* (x^2 + 2x +4) = x ^3 - 8 The 4 in the trinomial is multiplied by the -2 in the binomial to give -8. Your factorizatin will give x^3 - 16 when multiplied back out.

OpenStudy (anonymous):

Oh right! Because there's another -2 that's being multiplied by 4. Oh ok that makes sense.

Directrix (directrix):